The result of a survey shows that the amount of time a student spends on studying for a test is normally distributed. If the amount of time spent studying is 12 hours and the standard deviation is 4 hours, what is the probability that a student will spend more than 8 hours studying?


1
Expert's answer
2021-09-03T12:40:35-0400

"\mu=12 \\
\sigma=4 \\
P(X>8) = 1 -P(X≤8) \\
= 1 -P(Z≤ \frac{8-12}{4}) \\
= 1 -P(Z≤ -1) \\
= 1 -0.1586 \\
= 0.8414"